# UNO-Bench: A Unified Benchmark for Exploring the Compositional Law Between Uni-modal and Omni-modal in OmniModels

### v2

@@ Abstract
Multimodal Large Languages models have been progressing from uni-modal understanding toward unifying visual, audio and language modalities, collectively termed omni models. However, the correlation between uni-modal and omni-modal remains unclear, which requires comprehensive evaluation to drive omni model’s intelligence evolution. In this work, we propose a novel, high quality and UNified Omni model benchmark, UNO-Bench, which effectively assesses both UNi-modal and Omni-modal capabilities. The benchmark consists of 3730 human curated samples, with 98% cross-modality solvability, across 44 task types, and an innovative multi-step open-ended question type for assessing complex reasoning. Besides, a general scoring model supporting 6 question types is proposed for automated evaluation with 95% accuracy. Experimental result shows the Compositional Law between omni-modal and uni-modal performance and the omni-modal capability manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models. Our code and data are available at GitHub.
UNO-Bench: A Unified Benchmark for Exploring the Compositional Law Between Uni-modal and Omni-modal in OmniModels
Chen Chenthanks: Equal contribution.   Zeyang Hu11footnotemark: 1   Fengjiao Chenthanks: Project Lead. Liya Ma   Jiaxing Liu   Xiaoyu Li22footnotemark: 2   Xuezhi Cao Meituan, Beijing, China. {chenchen165, huzeyang, chenfengjiao02, maliya06}@meituan.com {liujiaxing10, lixiaoyu28, caoxuezhi}@meituan.com
@@ 1 Introduction
Existing omni model evaluation benchmarks remain relatively scarce and focus on different aspects. For example, some prioritize image comprehension(Li et al., 2024b), others emphasize video understanding(Hong et al., 2025), while a subset concentrates on speech interaction(Gong et al., 2024). Notably, existing datasets are exclusively English-centric, lacking evaluation benchmarks for Chinese linguistic contexts.
The ideal omni model should simultaneously preserve visual understanding capabilities (e.g., MMBench(Liu et al., 2024b)/MathVista(Lu et al., 2024a)/MVBench(Li et al., 2024a)), speech interaction proficiency (e.g., MMAU(Sakshi et al., 2025)), and cross-modal integration capacity (e.g., OmniBench(Li et al., 2024b)/WorldSense(Hong et al., 2025)). Current evaluation paradigms employ disjointed benchmark suites for separate capability testing, creating resource-intensive evaluation processes and disconnected modality assessments. Beyond uni-modal, omni-modal capability introduces advanced challenges across image, video and audio modality. However, 77% questions from WorldSense are solvable without vision or audio, and 25% questions from OmniBench contain erroneous answers. These issues limit the evaluation and analysis of omni models’ capabilities.
@@
Main Contributions:
2. Establish a high quality and diversity dataset construction pipeline including human-centric process and automated data compression. As as result, UNO-Bench comprises 1250 human curated samples for omni-modal with 98% cross-modality solvability, and 2480 enhanced samples for uni-modal, across 44 task types and 6 question types. The human-created novel dataset is well-suited to real-world scenarios, particularly within the Chinese context, whereas the automatically compressed dataset offers a 90% increase in speed and maintains 98% consistency across 18 public benchmarks. Its comprehensive quality and efficiency significantly surpasses existing datasets.
Figure 1: Illustration of the unified ability taxonomy proposed in UNO-Bench. (a) exemplarily presents the way to unify uni-modal and omni-modal samples. (b) shows the three-layer hierarchical structure of the unified ability taxonomy and its distribution in UNO-Bench. Among the 13 abilities, (c) demonstrates the $4$ diverse task types that cover a large range of real world scenarios.
@@ 2.1 Uni-Modal Benchmarks
Based on large language models, vision language models (VLM) (Bai et al., 2025; Xiaomi, 2025; Zeng et al., 2025) and audio language models (ALM) (Ding et al., 2025; Wu et al., 2025) introduce the general intelligence to vision modality and audio modality respectively. Various uni-modal benchmarks conduct comprehensive evaluations on VLMs (Liu et al., 2024b; Lu et al., 2024a; Wang et al., 2024b, c; Liu et al., 2024d; Mathew et al., 2021; Ouyang et al., 2024; Li et al., 2024a; Wu et al., 2024; Liu et al., 2024c; xAI, 2023; Xiao et al., 2021; Huang et al., 2025; Hu et al., 2025; Fu et al., 2024) and ALMs (Ardila et al., 2019; Wang et al., 2021; Yang et al., 2024; Ao et al., 2024). For image modality, MMBench(Liu et al., 2024b) proposed a systematically designed benchmark to evaluate general image understanding on 20 different tasks. Focused on mathematic, MathVision(Wang et al., 2024b) collected questions from 19 mathematic competitions to evaluate VLMs complex reasoning ability. In addition to above, OCRBench(Liu et al., 2024d) supplied the evaluation on text recognition and document understanding. For video modality, MVBench(Li et al., 2024a) aggregated 11 public video benchmarks and incorporated data enhancement process to cover 20 dynamic video understanding tasks. To complement the long video understanding area, LongVideoBench(Wu et al., 2024) introduces hourly video materials to evaluate the information retrieval ability from long context. For audio modality, MMAU(Sakshi et al., 2025) provides general audio understanding assessment across speech, sounds and music domains, featuring diverse audio samples. There’s massive uni-modal benchmarks covering diverse model abilities on vision modality and audio modality separately.
@@ 2.2 Omni-Modal Benchmarks
Omni models have arisen in recent years(Comanici et al., 2025; Xu et al., 2025b; AI et al., 2025; Li et al., 2025), as the pioneer, Gemini(Comanici et al., 2025) shows a strong ability in understanding both vision and audio, while Qwen-3-Omni(Xu et al., 2025b) provides leading performance in open-source models. However, there’s less omni-modal benchmarks that can evaluate the modality combination across image, video and audio. OmniBench(Li et al., 2024b) inserted audio as a context into the image understanding task and made up an omni-modal benchmark, while the data quality needs further improvement. WorldSence(Hong et al., 2025) emphasized audio-visual data in real world scenarios with high data quality, while most audio-visual questions can be solved by audio or video alone, which can’t assess the cross-modality ability. Other datasets focus on audio (Gong et al., 2024) or video (Zhou et al., 2025) and cover limited task types.
@@ 3 Method
In this section, we first introduce the omni-modal dataset construction pipeline in Section.3.1. The uni-modal dataset improvement method and a general evaluation dataset compression method to improve the evaluation efficiency are introduced in Section.3.2. Finally, the multi-step open-ended questions are introduced in Section.3.3.
@@ 3.1 Omni-modal Dataset Construction
Figure 2: Dataset Construction Pipeline includes human crafted process (left side) and automated data compression (right side). First, we collect diverse and novel materials to prevent data contamination. Second, with the proposed unified ability taxonomy, human annotators including experts will craft questions, answers and record audios in real-world scenarios. Finally, with model checking, ablation study and human experts revision, we achieves high quality and diversity dataset. Regarding automated data compression, we use model performances as the sample feature, employing clustering and resampling to achieve a 90% dataset compression.
@@ 3.1.1 Model Ability Taxonomy
Through cumulative experiences on multimodal evaluation from both model-side and user-side, we summarize the capabilities of uni-modal and omni-modal into a unified model ability taxonomy. As shown in Figure.1, the omni model’s capabilities are systematically categorized into two primary dimensions: Perception and Reasoning. Detailed definitions and examples can be found in the Appendix.E.
Perception dimension structured through seven recognition types: Object Perception, Attribute Perception, Scenario Perception, Spatial Perception, Cross-Modal Conversion, Semantic Understanding. In addition, we incorporate Cross-Modal Alignment evaluation to assess information synchronization across modalities.
The perception dimension fuses visual and audio modalities. For example, Scenario Perception includes the recognition of visual scenes and the judgment of audio scenes.
Reasoning dimension extends conventional reasoning categories (General, STEM, Code) with: Spatial Reasoning (including Static Reasoning and Dynamic Reasoning), Temporal Reasoning, and Complex Reasoning (multi-conditional, multi-step problem solving).
The unified ability taxonomy combines uni-modal and omni-modal abilities which provides a comprehensive measurement that is particularly critical for advanced omni models. Based on this taxonomy, we create a diversity dataset with 44 task types distributed as shown in Figure.1.
@@ 3.1.2 Material Collection
In both data quality checks and experimental results, we found that natural audio-visual synchronized video data contains a large amount of information redundancy, only a few videos require both audio and vision modality simultaneously. Therefore, we begin with carefully designed material collection. Our materials have the following three characteristics: a. Diverse Sources. The majority of our materials are real-world photos and videos collected through crowdsourcing, and another portion sourced from copyright-free website. Additionally, a small fraction comes from high-quality public datasets such as MMVU(Zhao et al., 2025), LongVideoBench(Wu et al., 2024), and VideoVista(Chen et al., 2025). b. Rich and Diverse Topics. Our materials cover a broad spectrum of subjects, including society, culture, art, life, literature, science, and so on. c. Live-Recorded Audio. Apart from background sounds and music, all dialogue is recorded by human speakers. With over 20 participants in the recording process, the audio features are rich and closely reflect the diverse vocal characteristics of the real world.
@@ 3.1.3 QA Annotation
Our annotators consist of human experts and high-quality crowd-sourced users. Human experts have extensive experience in cross-modal data construction and annotation, a deeper understanding of model capabilities, and thus ensure higher professionalism and specificity in the data they construct. Most crowd-sourced users are college students with rich experience in multi-modal model interactions and diverse professional backgrounds, providing data with better authenticity and diversity.
@@
Compared to conventional methods limited to human intervention only during the quality assurance phase, our pipeline integrates a "Human-in-the-Loop" approach, ensuring continuous manual involvement from the initial data sourcing to the final output. This methodology not only prevents data leakage but also more accurately simulates real-world scenarios. Furthermore, the manually curated Chinese dataset genuinely captures user requirements in a Chinese linguistic context, compensating for the shortcomings of most existing English-centric datasets.
@@ 3.1.4 Quality Inspection
To ensure the data quality, we have established a multi-stage, cyclically validated quality assurance system composed of automated tools and manual review. Each question undergoes at least three rounds of independent quality inspection to maximize data quality. First, a preliminary model check is conducted to filter out cases with ambiguous questions, non-unique answers, or those that do not conform to the task type. Second, through modality ablation experiments, we remove one modality of information from the QA pair to see if the model can answer based solely on the remaining information. If the question becomes unsolvable or ambiguous after removing any one modality, it proves the cross-modality solvability of the data. Finally, manual quality inspection and revision are performed.
@@ 3.2.1 Quality Improvement
Existing public uni-modal datasets are bothered by data leakage issue(Xu et al., 2024). To verify the influence, we adopt privatization improvement on the widely used public dataset MMBench(Liu et al., 2024b). As shown in Figure.3, the performance of models have better distinguishability after dataset improvement, reflecting the true capability differences between models. Therefore, for uni-modal data, we also follow the aforementioned construction process for self-construction datasets.
In addition to self-constructed data, we also selected some multimodal data from public datasets to supplement in terms of capability items and data types. (Data mainly comes from AV-Odyssey(Gong et al., 2024) and WorldSense(Hong et al., 2025), accounting for 11% of the total). The specific selection logic is as follows:
@@ 3.2.2 Dataset Compression
Regarding the existing large-scale uni-modal benchmarks, to reduce the evaluation cost of large-scale models, we designed a clustering-guided hierarchical sampling (CGHS) method as shown in Figure.2. This method achieves efficient compression and high evaluation consistency through the following steps:
@@
Cluster-based Stratification: Utilize the Kmeans++(Arthur and Vassilvitskii, 2007) algorithm to categorize questions into k clusters, each representing a "model performance similar" question type (e.g., easy questions, difficult questions, etc.).
@@
Validity Verification: Use Margin of Error (MoE) and Root Mean Square Error (RMSE) to measure theoretical and practical errors respectively. Verify the consistency between model rankings and scores using Spearman Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC).
@@ 3.3.1 Question Type Definition
Evaluating the multi-step reasoning capabilities of omni models presents a significant challenge. Real-world problems require models to integrate multi-modal information and execute a sequence of logical steps. However, current automated benchmarks, often relying on Outcome Reward Models (ORMs), typically provide only a binary pass/fail judgment. This approach fails to distinguish between a model that completes 80% of a task and one that fails at 20%, a crucial gap that human evaluators easily perceive. While alternatives like Process Reward Models (PRMs)Lightman et al. (2023) or multi-turn dialoguesReddy et al. (2019) exist, they are hampered by high implementation difficulty, low accuracy, or poor efficiency. Moreover, the prevalence of multiple-choice formats in existing benchmarks is unrepresentative of real-world, open-ended user queries and can mask model deficiencies.

### v3

@@ Abstract
Multimodal Large Languages models have been progressing from uni-modal understanding toward unifying visual, audio and language modalities, collectively termed omni models. However, the correlation between uni-modal and omni-modal remains unclear, which requires comprehensive evaluation to drive omni model’s intelligence evolution. In this work, we introduce a novel, high-quality, and UNified Omni model benchmark, UNO-Bench. This benchmark is designed to effectively evaluate both UNi-modal and Omni-modal capabilities under a unified ability taxonomy, spanning 44 task types and 5 modality combinations. It includes 1250 human curated samples for omni-modal with 98% cross-modality solvability, and 2480 enhanced uni-modal samples. The human-generated dataset is well-suited to real-world scenarios, particularly within the Chinese context, whereas the automatically compressed dataset offers a 90% increase in speed and maintains 98% consistency across 18 public benchmarks. In addition to traditional multi-choice questions, we propose an innovative multi-step open-ended question format to assess complex reasoning. A general scoring model is incorporated, supporting 6 question types for automated evaluation with 95% accuracy. Experimental result shows the Compositional Law between omni-modal and uni-modal performance and the omni-modal capability manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
GitHub: https://github.com/meituan-longcat/UNO-Bench
Hugging Face: https://huggingface.co/datasets/meituan-longcat/UNO-Bench
Figure 1: Benchmark Statistics and Evaluation Results.
@@ 1 Introduction
Existing omni model evaluation benchmarks remain relatively scarce and focus on different aspects. For example, some prioritize image comprehension(Li et al., 2024a), others emphasize video understanding(Hong et al., 2025), while a subset concentrates on speech interaction(Gong et al., 2024). Notably, existing datasets are exclusively English-centric, lacking evaluation benchmarks for Chinese linguistic contexts.
The ideal omni model should simultaneously preserve visual understanding capabilities (e.g., MMBench(Liu et al., 2024a)/MathVista(Lu et al., 2024a)/MVBench(Li et al., 2024b)), speech interaction proficiency (e.g., MMAU(Sakshi et al., 2025)), and cross-modal integration capacity (e.g., OmniBench(Li et al., 2024a)/WorldSense(Hong et al., 2025)). Current evaluation paradigms employ disjointed benchmark suites for separate capability testing, creating resource-intensive evaluation processes and disconnected modality assessments. Beyond uni-modal, omni-modal capability introduces advanced challenges across image, video and audio modality. However, 77% questions from WorldSense are solvable without vision or audio, and 25% questions from OmniBench contain erroneous answers. These issues limit the evaluation and analysis of omni models’ capabilities.
Main Contributions:
2. Establish a high quality and diversity dataset construction pipeline including human-centric process and automated data compression. As a result, UNO-Bench comprises 1250 human curated samples for omni-modal with 98% cross-modality solvability, and 2480 enhanced samples for uni-modal, across 44 task types and 5 modality combinations. The human-created novel dataset is well-suited to real-world scenarios, particularly within the Chinese context, whereas the automatically compressed dataset offers a 90% increase in speed and maintains 98% consistency across 18 public benchmarks. Its comprehensive quality and efficiency significantly surpasses existing datasets.
@@ 2.1 Uni-Modal Benchmarks
Based on large language models, vision language models (VLM) (Bai et al., 2025; Xiaomi, 2025; Zeng et al., 2025) and audio language models (ALM) (Ding et al., 2025; Wu et al., 2025) introduce the general intelligence to vision modality and audio modality respectively. Various uni-modal benchmarks conduct comprehensive evaluations on VLMs (Liu et al., 2024a; Lu et al., 2024a; Wang et al., 2024a, b; Liu et al., 2024b; Mathew et al., 2021; Ouyang et al., 2024; Li et al., 2024b; Wu et al., 2024; Liu et al., 2024c; xAI, 2023; Xiao et al., 2021; Huang et al., 2025; Hu et al., 2025; Fu et al., 2024) and ALMs (Ardila et al., 2019; Wang et al., 2021; Yang et al., 2024; Ao et al., 2024). For image modality, MMBench(Liu et al., 2024a) proposed a systematically designed benchmark to evaluate general image understanding on 20 different tasks. Focused on mathematic, MathVision(Wang et al., 2024a) collected questions from 19 mathematic competitions to evaluate VLMs complex reasoning ability. In addition to above, OCRBench(Liu et al., 2024b) supplied the evaluation on text recognition and document understanding. For video modality, MVBench(Li et al., 2024b) aggregated 11 public video benchmarks and incorporated data enhancement process to cover 20 dynamic video understanding tasks. To complement the long video understanding area, LongVideoBench(Wu et al., 2024) introduces hourly video materials to evaluate the information retrieval ability from long context. For audio modality, MMAU(Sakshi et al., 2025) provides general audio understanding assessment across speech, sounds and music domains, featuring diverse audio samples. There are massive uni-modal benchmarks covering diverse model abilities on vision modality and audio modality separately.
@@ 2.2 Omni-Modal Benchmarks
Omni models have arisen in recent years(Comanici et al., 2025; Xu et al., 2025a; AI et al., 2025; Li et al., 2025), as the pioneer, Gemini(Comanici et al., 2025) shows a strong ability in understanding both vision and audio, while Qwen-3-Omni(Xu et al., 2025a) provides leading performance in open-source models. However, there are less omni-modal benchmarks that can evaluate the modality combination across image, video and audio. OmniBench(Li et al., 2024a) inserted audio as a context into the image understanding task and made up an omni-modal benchmark, while the data quality needs further improvement. WorldSence(Hong et al., 2025) emphasized audio-visual data in real world scenarios with high data quality, while most audio-visual questions can be solved by audio or video alone, which cannot assess the cross-modality ability. Other datasets focus on audio (Gong et al., 2024) or video (Zhou et al., 2025) and cover limited task types. For instance, in Figure.3(b), the problem can be resolved using either the audio modality or the visual modality, whereas in Figure.3(c), only the visual modality is necessary to address the problem. These instances are likely to exaggerate the capabilities of the omni model, making it crucial to evaluate the cross-modality solvable problem (illustrated in Figure.3(a)) to accurately assess omni-modal capability (refer to Section.4.3 for more details).
@@ 3 Method
In this section, we first introduce the omni-modal dataset construction pipeline in Section.3.1. For uni-modal dataset, an quality improvement method and a general dataset compression method to improve the evaluation efficiency are introduced in Section.3.2. Finally, the novel multi-step open-ended questions are presented alongside a general scoring model in Section.3.3
@@ 3.1 Omni-modal Dataset Construction
Figure 2: Illustration of the unified ability taxonomy proposed in UNO-Bench.
@@ 3.1.1 Model Ability Taxonomy
Through cumulative experiences on multimodal evaluation from both model-side and user-side, we summarize the capabilities of uni-modal and omni-modal into a unified model ability taxonomy. As shown in Figure.2(b), the omni model’s capabilities are systematically categorized into two primary dimensions: Perception and Reasoning. Detailed definitions and examples can be found in the Appendix.D.
Perception dimension structured through seven recognition types including Object Perception, Attribute Perception, Scenario Perception, Spatial Perception, Cross-Modal Conversion, Semantic Understanding. In addition, we incorporate Cross-Modal Alignment to assess information synchronization across modalities.
Reasoning dimension extends conventional reasoning categories (including General, STEM, Code) with Spatial Reasoning (including Static Reasoning and Dynamic Reasoning), Temporal Reasoning, and Complex Reasoning (which indicates multi-conditional, multi-step problem).
As shown in Figure.2(a), the unified ability taxonomy combines uni-modal and omni-modal abilities which provides a comprehensive measurement that is particularly critical for omni models. For example, Scenario Perception includes the recognition of visual scenes and the judgment of audio scenes. Based on this taxonomy, we create a diversity dataset with 44 task types illustrated in Figure.2(c).
@@ 3.1.2 Material Collection
In both data quality checks and experimental results, we found that the natural video with audio-visual synchronized data contains a large amount of information redundancy, only a few videos require both audio and visual modality simultaneously. Therefore, we begin with carefully designed material collection.
Our materials have the following three characteristics:
Diverse Sources. The majority of our materials are real-world photos and videos collected through crowd sourcing, and another portion sourced from copyright-free websites. Additionally, a small fraction comes from high-quality public datasets such as MMVU(Zhao et al., 2025), LongVideoBench(Wu et al., 2024), and VideoVista(Chen et al., 2025).
Rich and Diverse Topics. Our materials cover a broad spectrum of subjects, including society, culture, art, life, literature, science, and so on.
Live-Recorded Audio. Apart from background sounds and music, all dialogue is recorded by human speakers. With over 20 participants in the recording process, the audio features are rich and closely reflect the diverse vocal characteristics of the real world, such as Mandarin and Sichuan dialect.
@@
Figure 3: Illustration of the cross-modality solvable sample.
@@ 3.1.3 QA Annotation
Our annotators consist of human experts and high-quality crowd-sourced users. Human experts have extensive experience in cross-modal data construction and annotation, a deeper understanding of model capabilities, and thus ensure higher professionalism and specificity in the data they construct. Most crowd-sourced users are college students with rich experience in multimodal model interactions and diverse professional backgrounds, providing data with better authenticity and diversity.
@@
Compared to conventional methods limited to human intervention only during the quality assurance phase, our pipeline integrates a human-centric approach, ensuring continuous manual involvement from the initial data sourcing to the final output. This methodology not only prevents data leakage but also more accurately simulates real-world scenarios. Furthermore, the manually curated Chinese dataset genuinely captures user requirements in a Chinese linguistic context, compensating for the shortcomings of most existing English-centric datasets.
@@ 3.1.4 Quality Inspection
To ensure the data quality, we have established a multi-stage, cyclically validated quality assurance system composed of automated tools and manual review. Each question undergoes at least three rounds of independent quality inspection to maximize data quality. Model Check, a preliminary model check is conducted to filter out cases with ambiguous questions, non-unique answers, or those that do not conform to the task type. Ablation Study, through modality ablation experiments, we remove one modality of information from the QA pair to see if the model can answer based solely on the remaining information. If the question becomes unsolvable or ambiguous after removing any one modality, it proves the cross-modality solvability of the data. Human Check, finally manual quality inspection and revision are performed.
Figure 4: Dataset Construction Pipeline includes human-centric process (left side) and automated data compression (right side). First, we collect diverse and novel materials to prevent data contamination. Second, with the proposed unified ability taxonomy, human annotators including experts will craft questions, answers and record audios in real-world scenarios. Finally, with model checking, ablation study and human experts revision, we achieves high quality and diversity dataset. Regarding automated data compression, we present a clustering-guided hierarchical sampling method to achieve efficient compression while maintaining high evaluation consistency.
@@ 3.2.1 Quality Improvement
Existing public uni-modal datasets are bothered by data leakage issue(Xu et al., 2024). To verify the influence, we adopt privatization improvement on the widely used public dataset MMBench(Liu et al., 2024a). As shown in Figure.13, the performance of models have better distinguishability after dataset improvement, reflecting the true capability differences between models. Therefore, for uni-modal data, we also follow the aforementioned construction process for self-construction datasets. In addition to self-constructed data, we also selected some multimodal data from public datasets to supplement in terms of capability items and data types. (Data mainly comes from AV-Odyssey(Gong et al., 2024) and WorldSense(Hong et al., 2025), accounting for 11% of the total). The specific selection logic is as follows:
@@ 3.2.2 Dataset Compression
Regarding the existing large-scale uni-modal benchmarks, to reduce the evaluation cost of large-scale models, we designed a clustering-guided hierarchical sampling (CGHS) method as shown in Figure.4. CGHS is a general method for dataset compression, which utilizes model performance metrics as features rather than the content of questions to select important samples that impact model performance. For training datasets, CGHS can retain both simple and difficult samples in an unsupervised manner or minimize similar rollout samples in a batch for online policy. When it comes to test datasets, CGHS is capable of achieving efficient compression while maintaining high evaluation consistency. The introduction of CGHS is outlined in the following steps:
@@
Cluster-based Stratification: Utilize the Kmeans++(Arthur and Vassilvitskii, 2007) algorithm to categorize questions into k clusters, each representing a ”model performance similar” question type (e.g., easy questions, difficult questions, etc.).
@@
Validity Verification: To verify the compression performance, we define these metrics: Spearman’s Rank Correlation Coefficient (SRCC) for ranking consistency, Pearson’s Linear Correlation Coefficient (PLCC) for linear value consistency, Root Mean Square Error (RMSE) for numerical precision, Margin of Error (MoE) for quantifying estimation uncertainty, and Confidence Interval Coverage (CIC) for statistical reliability.
@@ 3.3.1 Question Type Definition
Evaluating the multi-step reasoning capabilities of omni models presents a significant challenge. Real-world problems require models to integrate multi-modal information and execute a sequence of logical steps. However, current automated benchmarks, often relying on Outcome Reward Models (ORMs), typically provide only a binary pass/fail judgment. This approach fails to distinguish between a model that completes 80% of a task and one that fails at 20%, a crucial gap that human evaluators easily perceive. While alternatives like Process Reward Models (PRMs)Lightman et al. (2023) or multi-turn dialoguesReddy et al. (2019) exist, they are hampered by high implementation difficulty, low accuracy, or poor efficiency. Moreover, the prevalence of multiple-choice formats in existing benchmarks is unrepresentative of real-world, open-ended user queries and may conceal the weaknesses of models.